OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). 6x – 2y = 2 –3x + 4y = 5 A. (1, 3) B. (2, 3) C. (1, –2) D. (1, 2)

OpenStudy (jjuden):

its d

OpenStudy (jjuden):

-3(1)+4(2)=5 -3+8=5 6(1)-2(2)=2 6-4=2